OpenStudy (anonymous):

What is the length of side x? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

posting a pic

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah i was going to say that too

OpenStudy (anonymous):

so that would be

OpenStudy (igreen):

Pythagorean theorem. \(a^2 + b^2 = c^2\) Where 'a' and 'b' are the two legs, and 'c' is the hypotenuse. We are given a leg and a hypotenuse. Plug them in: \(23^2 + b^2 = 31^2\)

OpenStudy (igreen):

Solve for 'b'

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@ 23^+B^=31^

OpenStudy (anonymous):

haha didnt see you already wrote it out for me

OpenStudy (igreen):

Simplify 23^2 and 31^2. Can you do that?

OpenStudy (anonymous):

Yeah Dude

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

529 and 961

OpenStudy (igreen):

Yes, so we have: 529 + b^2 = 961 Now subtract 529 to both sies.

OpenStudy (igreen):

*sides

OpenStudy (igreen):

b^2 = 961 - 529 What's 961 - 529?

OpenStudy (anonymous):

sorry using B-room

OpenStudy (anonymous):

$32

OpenStudy (anonymous):

432*

OpenStudy (igreen):

Yes, so we have: \(b^2 = 432\) Now we find the square root of both sides: \(b = \sqrt{432}\) Can you find the square root of 432?

OpenStudy (anonymous):

20.78

OpenStudy (igreen):

Correct! That's our answer :)

OpenStudy (anonymous):

Thank Again green